What is a ratio? Give several examples of ratios in everyday use.


1
Expert's answer
2022-02-07T15:27:54-0500

A ratio is a fraction that shows how one value relates to other, or one value divided by other. For instance, we have two cars: one moving at 20 km/h and the other at 60 km/h. The ratio of their velocities is 20/60=1/3.

The temperature in one vessel is 50 K, and 15 K in other. The ratio is 50/15=10/3.